OpenStudy (anonymous):

What is the length of stack A G with bar on top in the rectangular prism? Round your answer to the nearest tenth. A. 12.6 cm B. 13.0 cm C. 13.6 cm D. 14.2 cm

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Do you know the Pythagorean Theorem?

OpenStudy (anonymous):

A^2+B^2+C^2

OpenStudy (joannablackwelder):

Close, a^2+b^2=c^2

OpenStudy (joannablackwelder):

So, can you use that to find DG?

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

Cool, what do you get?

OpenStudy (anonymous):

169

OpenStudy (joannablackwelder):

That is c^2. What is c?

OpenStudy (anonymous):

144

OpenStudy (joannablackwelder):

Wait, so

OpenStudy (joannablackwelder):

You got 169 the first time. Which is what I get for 5^2 + 12^2

OpenStudy (joannablackwelder):

And that equals 169 which is equal to c^2

OpenStudy (joannablackwelder):

169=c^2

OpenStudy (joannablackwelder):

Can you solve for c?

OpenStudy (anonymous):

wait so add 169 & c^2

OpenStudy (joannablackwelder):

No,  So, we need to find a number that when we square it, it equals 169

OpenStudy (anonymous):

when i do that is says error

OpenStudy (joannablackwelder):

Hm, what do you mean?

OpenStudy (anonymous):

13

OpenStudy (joannablackwelder):

Right :-)

OpenStudy (joannablackwelder):

So, now we are looking for AG. Any thoughts on how to get that?

OpenStudy (anonymous):

same way ?

OpenStudy (joannablackwelder):

Yep. Using what right triangle?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

No, that's not what I get.

OpenStudy (joannablackwelder):

What triangle are you looking at?

OpenStudy (anonymous):

AG

OpenStudy (joannablackwelder):

AG isn't a triangle. It's a line.

OpenStudy (anonymous):

C

OpenStudy (joannablackwelder):

How did you get that?

OpenStudy (anonymous):

Do I Need To Be Looking For Angle or did i already get the answer

OpenStudy (joannablackwelder):

You are looking for the length of AG.